# Quantum Computation by Cooling

### v2

@@ V Discussion
In fact, the idea of CQC can be used to prepare the ground states of general quantum Hamiltonians, although we focused on the analyzable cases. The most essential requirement is that the ground state should not be heated to make it a steady state. This is ensured when the ground state has a finite spectral gap to the excited state and this spectral gap is resolved. Consequently, the inverse of the spectral gap determines the basic time scale of the cooling, which is a natural consequence of the energy-time uncertainty principle. For gapped Hamiltonians, as the transition strength can be determined finitely without the risk of heating the ground state, a finite number of cavity modes can cover the transitions across the entire spectrum in principle, where the number is roughly given by the ratio of the whole energy range to the characteristic transition rate. This means that the number of cavity modes increases only polynomially with the system size for gapped Hamiltonians. However, choosing an appropriate transition term is heavily dependent on the Hamiltonian, hence analytically challenging.

### v4

@@ V Discussion
The idea of CQC could be used to prepare the ground states of general gapped quantum Hamiltonians. A qualitative argument is the following. Recall that the transition from $|z_{1}\rangle_{P}$ to $|z_{2}\rangle_{P}$ occurs when the condition $|E(z_{2})-\{E(z_{1})-\omega_{m}\}|\lesssim|{}_{P}\langle z_{2}|\lambda H_{T}|z_{1}\rangle_{P}|$ is satisfied, where a cavity mode with frequency $\omega_{m}$ is involved. The key point is that a single cavity mode enables transitions with the associated energy difference up to a margin determined by the spectral linewidth $|{}_{P}\langle z_{2}|\lambda H_{T}|z_{1}\rangle_{P}|$. To ensure that the transition can distinguish the ground state from the excited state, this linewidth should be sufficiently smaller than the spectral gap above the ground state. Consequently, a finite spectral gap sets the upper bound on the transition rate, hence the time scale of the cooling. Crucially, a finite spectral gap in turn allows to have a finite linewidth for transitions. The entire spectral range of the problem Hamiltonian, which is $\mathcal{O}(N)$, can then be divided into $\mathcal{O}(N/\lambda)$ sectors. This means that a polynomial number of cavity modes suffice to cover the entire spectrum, provided that one can design appropriate transition terms $H_{T}$ along with the cavity modes to establish transition channels to the ground state for all the energy levels. Of course, the practical application of this idea would heavily depend on the specific Hamiltonian and hence analytically challenging. The requirement of a finite spectral gap is a natural consequence of the energy-time uncertainty principle.